4x+9 For x=8, -4x^2+2x+7 for x=-1, a^2+5abc-c^2 for a=-8,b=3, and c=5

Answer by checkley71(8403)

4X+9 FOR X=8
4*8+9=32+9=41 ANSWER.
---------------------------------------
-4X^2+2X+7 FOR X=-1
-4*-1^2+2*-1+7=-4*1-2+7=-4-2+7=1 ANSWER.
-------------------------------------------
A^2+5ABC-C^2 FOR A=-8, B=3, C=5
-8^2+5*-8*3*5-5^2=64-600+25=-511 ANSWER.